# The N3LO Twist-2 Matching of Helicity TMDs and SIDIS $q_{\ast}$ Spectrum

### v3

@@ 4.5 N3LO coefficient functions and small-x expansion
$\Delta\mathscr{I}^{\text{s}(1)}_{qq}(x)=$ $2C_{F}\,$, (150)
$\Delta\mathscr{I}^{\text{s}(2)}_{qq}(x)=$ $C_{A}C_{F}\left(\left(-\frac{2}{3}\right)\ln^{3}x+\left(-\frac{11}{6}\right)\ln^{2}x+\left(-\frac{31}{9}\right)\ln x-12\zeta_{2}+4\zeta_{3}+\frac{397}{27}\right)$
@@
$+$ $C_{F}^{2}\left(\ln^{3}x+3\ln^{2}x+4\ln x+18\zeta_{2}+20\zeta_{3}-52\right)\,$, (151)
$\Delta\mathscr{I}^{\text{s}(1)}_{qg}(x)=$ $8N_{f}T_{F}\,$, (152)
$\Delta\mathscr{I}^{\text{s}(2)}_{qg}(x)=$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+22\ln^{2}x+72\ln x-32\zeta_{2}+8\zeta_{3}+140\right)$
@@
$\Delta\mathscr{I}^{\text{s}(3)}_{qg}(x)=$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(\frac{32308}{81}-12\zeta_{2}\right)\ln^{3}x+\left(\left(-\frac{424}{3}\right)\zeta_{2}+\left(-\frac{328}{3}\right)\zeta_{3}+\frac{19276}{9}\right)\ln^{2}x$
@@
$+$ $\left(-\frac{764}{27}\right)\ln^{4}x+\left(-\frac{23300}{81}\right)\ln^{3}x+\left(-\frac{848}{9}\right)\zeta_{4}+\frac{2416}{81}\zeta_{2}+\frac{5984}{27}\zeta_{3}-\frac{423361}{81}\bigg]\,$, (154)
$\Delta\mathscr{I}^{\text{s}(1)}_{gq}(x)=$ $-4C_{F}\,$, (155)
$\Delta\mathscr{I}^{\text{s}(2)}_{gq}(x)=$ $C_{A}C_{F}\left(\left(-\frac{4}{3}\right)\ln^{3}x-16\ln^{2}x-50\ln x+28\zeta_{2}+48\zeta_{3}-\frac{4454}{27}\right)$
@@
$\Delta\mathscr{I}^{\text{s}(3)}_{gq}(x)=$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\frac{64}{9}\zeta_{2}+\frac{19232}{81}\right)\ln^{2}x+\left(\frac{3904}{27}\zeta_{2}+\frac{1472}{9}\zeta_{3}+\frac{110612}{243}\right)\ln x$
@@
$+$ $\zeta_{2}\left(144\zeta_{3}+284\right)-640\zeta_{5}-\frac{12365}{12}\bigg]\,$, (157)
$\Delta\mathscr{I}^{\text{s}(1)}_{gg}(x)=$ $-8C_{A}\,$, (158)
$\Delta\mathscr{I}^{\text{s}(2)}_{gg}(x)=$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{2}x+\frac{136}{9}\ln x+\frac{1882}{27}\right)+C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+18\ln^{2}x+84\ln x+120\right)$
@@
$\Delta\mathscr{I}^{\text{s}(3)}_{gg}(x)=$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(8\zeta_{2}+\frac{3416}{81}\right)\ln^{3}x+\left(\frac{488}{9}\zeta_{2}+112\zeta_{3}+\frac{10708}{27}\right)\ln^{2}x$

### v4

@@ 4.5 N3LO coefficient functions and small-x expansion
$\Delta\mathscr{I}^{\text{s}(1)}_{qq}(x)\simeq$ $2C_{F}\,$, (141)
$\Delta\mathscr{I}^{\text{s}(2)}_{qq}(x)\simeq$ $C_{A}C_{F}\left(\left(-\frac{2}{3}\right)\ln^{3}x+\left(-\frac{11}{6}\right)\ln^{2}x+\left(-\frac{31}{9}\right)\ln x-12\zeta_{2}+4\zeta_{3}+\frac{397}{27}\right)$
@@
$+$ $C_{F}^{2}\left(\ln^{3}x+3\ln^{2}x+4\ln x+18\zeta_{2}+20\zeta_{3}-52\right)\,.$ (142)
$\Delta\mathscr{I}^{\text{s}(1)}_{qg}(x)\simeq$ $8N_{f}T_{F}\,$, (143)
$\Delta\mathscr{I}^{\text{s}(2)}_{qg}(x)\simeq$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+22\ln^{2}x+72\ln x-32\zeta_{2}+8\zeta_{3}+140\right)$
@@
$\Delta\mathscr{I}^{\text{s}(3)}_{qg}(x)\simeq$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(\frac{32308}{81}-12\zeta_{2}\right)\ln^{3}x+\left(\left(-\frac{424}{3}\right)\zeta_{2}+\left(-\frac{328}{3}\right)\zeta_{3}+\frac{19276}{9}\right)\ln^{2}x$
@@
$+$ $\left(-\frac{764}{27}\right)\ln^{4}x+\left(-\frac{23300}{81}\right)\ln^{3}x+\left(-\frac{848}{9}\right)\zeta_{4}+\frac{2416}{81}\zeta_{2}+\frac{5984}{27}\zeta_{3}-\frac{423361}{81}\bigg]\,.$ (145)
$\Delta\mathscr{I}^{\text{s}(1)}_{gq}(x)\simeq$ $-4C_{F}\,$, (146)
$\Delta\mathscr{I}^{\text{s}(2)}_{gq}(x)\simeq$ $C_{A}C_{F}\left(\left(-\frac{4}{3}\right)\ln^{3}x-16\ln^{2}x-50\ln x+28\zeta_{2}+48\zeta_{3}-\frac{4454}{27}\right)$
@@
$\Delta\mathscr{I}^{\text{s}(3)}_{gq}(x)\simeq$ $C_{A}C_{F}N_{f}T_{F}\bigg[\left(\frac{64}{9}\zeta_{2}+\frac{19232}{81}\right)\ln^{2}x+\left(\frac{3904}{27}\zeta_{2}+\frac{1472}{9}\zeta_{3}+\frac{110612}{243}\right)\ln x$
@@
$+$ $\zeta_{2}\left(144\zeta_{3}+284\right)-640\zeta_{5}-\frac{12365}{12}\bigg]\,.$ (148)
$\Delta\mathscr{I}^{\text{s}(1)}_{gg}(x)\simeq$ $-8C_{A}\,$, (149)
$\Delta\mathscr{I}^{\text{s}(2)}_{gg}(x)\simeq$ $C_{A}N_{f}T_{F}\left(\frac{4}{3}\ln^{2}x+\frac{136}{9}\ln x+\frac{1882}{27}\right)+C_{F}N_{f}T_{F}\left(\frac{4}{3}\ln^{3}x+18\ln^{2}x+84\ln x+120\right)$
@@
$\Delta\mathscr{I}^{\text{s}(3)}_{gg}(x)\simeq$ $C_{A}^{2}N_{f}T_{F}\bigg[\left(8\zeta_{2}+\frac{3416}{81}\right)\ln^{3}x+\left(\frac{488}{9}\zeta_{2}+112\zeta_{3}+\frac{10708}{27}\right)\ln^{2}x$